OpenStudy (anonymous):

What is an equation in slope-intercept form for the line that passes through the points (1, –3) and (3, 1)? y = 3x + 1 y = x – 3 y = 2x + 5 y = 2x – 5

OpenStudy (campbell_st):

can you find the slope of the line...?

OpenStudy (anonymous):

yes its -2

OpenStudy (anonymous):

well then just plot one of your points and keep using your slope.. down two, right one, point. down two right one point. or vice versa up two, left one, point

OpenStudy (anonymous):

then use slope int. form y=mx+b where m=slope and b=y-int.

OpenStudy (campbell_st):

ok, so the slope intercept form of a line is \[y = mx + b\] and using your slope you have y = -2x + b so to find b pick a point (3, 1) and substitute x = 3 and y = 1 to find b. then you'll have the equation of the line.

OpenStudy (anonymous):

so b=7?

OpenStudy (campbell_st):

but looking at your choices... your slope calculation is incorrect... (1 - -3)/(3 - 1)

OpenStudy (anonymous):

ohh...

OpenStudy (campbell_st):

lets get the slope right 1st

OpenStudy (anonymous):

I still got -2..

OpenStudy (anonymous):

1- (-3) = -4. 3-1= 2. so thats -4/2, simplified to -2/1= -2

OpenStudy (campbell_st):

1 - -3 = 4 2 negatives make a positive so the slope is 4/2 or 2 so you have y = 2x + b substitute x = 3, y = 1 to find b and the equation of your line

OpenStudy (anonymous):

so b= -5?

OpenStudy (campbell_st):

thats it so your line is y = 2x - 5

OpenStudy (anonymous):

omg thank you:D <3333

OpenStudy (campbell_st):

glad to help